# GeneMAN: Generalizable Single-Image 3D Human Reconstruction from Multi-Source Human Data

### v2

@@ Abstract
Given a single in-the-wild human photo, it remains a challenging task to reconstruct a high-fidelity 3D human model. Existing methods face difficulties including a) the varying body proportions captured by in-the-wild human images; b) diverse personal belongings within the shot; and c) ambiguities in human postures and inconsistency in human textures. In addition, the scarcity of high-quality human data intensifies the challenge. To address these problems, we propose a Generalizable image-to-3D huMAN reconstruction framework, dubbed GeneMAN, building upon a comprehensive multi-source collection of high-quality human data, including 3D scans, multi-view videos, single photos, and our generated synthetic human data. GeneMAN encompasses three key modules. 1) Without relying on parametric human models (e.g., SMPL), GeneMAN first trains a human-specific text-to-image diffusion model and a view-conditioned diffusion model, serving as GeneMAN 2D human prior and 3D human prior for reconstruction, respectively. 2) With the help of the pretrained human prior models, the Geometry Initialization-&\&&-Sculpting pipeline is leveraged to recover high-quality 3D human geometry given a single image. 3) To achieve high-fidelity 3D human textures, GeneMAN employs the Multi-Space Texture Refinement pipeline, consecutively refining textures in the latent and the pixel spaces. Extensive experimental results demonstrate that GeneMAN could generate high-quality 3D human models from a single image input, outperforming prior state-of-the-art methods. Notably, GeneMAN could reveal much better generalizability in dealing with in-the-wild images, often yielding high-quality 3D human models in natural poses with common items, regardless of the body proportions in the input images.
@@ 1 Introduction
Creating high-quality 3D human models is crucial in various real applications, including VR/AR, telepresence, digital human interfaces, film, and 3D game production. Traditional methods [5, 24, 62] usually utilize a dense camera array to capture synchronized posed multi-view images for human reconstruction, which typically involves complicated and time-consuming processes. Towards efficient 3D human reconstruction, many approaches [47, 73, 48, 18, 61, 72, 13] delve into the challenge of reconstructing 3D human models from a single image, which, however, remains an ill-posed problem due to the absence of comprehensive 3D human observation data.
@@
To facilitate 3D human reconstruction, previous methods [73, 63, 64, 18, 72, 13] often employ parametric human models, e.g., SMPL [30] or SMPL-X [38] as 3D human geometry prior. Nevertheless, these parametric models fail to capture 3D clothing details, which impedes the accurate reconstruction of 3D human figures, particularly when individuals are adorned in loose garments. Recently, image-to-3D human reconstruction methods [13, 18, 72] integrate pretrained text-to-image diffusion models [44, 46] to incorporate 2D priors into the reconstruction process. Despite advances in generalizability, single-view 3D human reconstruction has not been fully resolved, especially when dealing with in-the-wild images. The primary challenges are shown in Fig. 1: (a) Varying Body Proportions: many portraits are captured with varying body proportions, such as full-body, half-body, and headshots, while existing methods primarily focus on full-body reconstructions; (b) Human with Objects: in everyday photography, it is common to capture people holding objects, standing on items, or wearing various accessories, which could greatly impact the reconstruction quality. (c) Human Reconstruction with Natural Pose and Textures: due to the absence of a broadly applicable human-specific geometry and texture model, existing methods struggle to reconstruct credible geometry and consistent texture from real-world images. Additionally, the scarcity of high-quality human body data exacerbates the difficulty of tackling the problem.
@@
In this paper, we propose GeneMAN, a Generalizable image-to-3D huMAN reconstruction framework for high-fidelity 3D human reconstruction from a single in-the-wild image. To enhance generalizability, we first collect a comprehensive, multi-source training dataset of high-quality, multi-modal human data, including 3D scans, multi-view videos, single images, and our augmented human data. Based on the multi-source human data collection, human-specific prior models, including a text-to-image diffusion model as the 2D prior and a view-conditioned diffusion model as the 3D prior, have been trained, which could provide more generalizable human priors compared to traditional human parametric models. Leveraging the pretrained human prior models, GeneMAN generates 3D human models through two primary stages. 1) 3D Geometry Reconstruction. To create precise and intricate 3D human geometry, GeneMAN employs an Initialization-&\&&-Sculpting strategy, which initially predicts a coarse human geometry using NeRF [36], followed by a sculpting process for adding geometry details. 2) 3D Texture Generation. Utilizing the refined human geometry, GeneMAN implements a multi-space texture refinement pipeline to produce high-fidelity, consistent 3D textures. Initially, coarse textures are created through multi-view texturing and iteratively refined in the latent space. Subsequently, detailed 3D textures are achieved via pixel space texture refinement, optimizing the UV maps with a 2D prior-based ControlNet [68]. Extensive experiments demonstrate that GeneMAN surpasses existing state-of-the-art (SoTA) methods, showcasing strong generalization capability and high generation quality. We would like to highlight that GeneMAN could faithfully reconstruct 3D human models with diverse clothing, complex poses, and different personal belongings, given a single in-the-wild image with varying body proportions. Our contributions are summarized as follows:
@@
  • A few effective 3D human reconstruction modules, such as Geometry Initialization &\&& Sculpting, and Multi-Space Texture Refinement, have been proposed, facilitating template-free 3D human geometry modeling and view-consistency texturing.
@@ 2 Related Work
Template-based 3D Human Reconstruction. Template-based approaches rely on human parametric models [30, 39, 65] to guide 3D reconstruction [73, 63, 64, 19, 11, 71, 72, 13]. PaMIR [73] and ICON [63] leverage SMPL-derived features to facilitate implicit surface regression. By contrast, ECON [64] employs explicit representations, predicting normals and depths to generate 2.5D surfaces. GTA [71] utilizes transformers with learnable embeddings to translate image features into 3D triplane features. Several methods [18, 13, 72] exploit the generative capabilities of pretrained diffusion models [44] for 3D human reconstruction. SiTH [13] finetunes an image-conditioned diffusion model to hallucinate back views, while SIFU [72] and TeCH [18] utilize diffusion-based geometry and texture refinement to improve reconstruction quality. Recently, IDOL [74] predicts Gaussian-based representation of 3D human to enable fast reconstruction. However, template-based methods are highly reliant on precise human parametric models, which are frequently unattainable for in-the-wild human images.
@@
Template-Free 3D Human Reconstruction. Without parameterized human models, PIFu [47] introduces a novel approach by extracting pixel-aligned features to build neural fields. PIFuHD [48] improves this with high-resolution normal guidance, while PHOHRUM [4] jointly predicts 3D geometry, surface albedo, and shading. Recently, HumanSGD [3] leverages pretrained 2D diffusion models [44] as a human appearance prior. HumanLRM [61] trains a diffusion-guided feed-forward model to predict the implicit field of a human without template priors. Nonetheless, they still encounter challenges, such as unrealistic and inconsistent textures, and inferior geometric details, largely due to insufficient priors and overly simplistic model designs.
@@ 3 Preliminaries
Diffusion Model. Diffusion models [14, 55, 56] learn to transform samples from a tractable noise distribution towards a data distribution. They comprise a forward process $\{q_{t}\}_{t\in[0,1]}$, which progressively adds random noise to a data point $\bm{x}_{0}\sim q_{0}(\bm{x}_{0})$, and a reverse process $\{p_{t}\}_{t\in[0,1]}$, which gradually recovers clean data from noise. The forward process is defined by a conditional Gaussian distribution: $q_{t}(\bm{x}_{t}|\bm{x}_{0}):=\mathcal{N}(\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$, where $\alpha_{t},\sigma_{t}>0$ are time-dependent coefficients. The reverse process is defined by denoising from $p_{1}(\bm{x}_{1})\coloneqq\mathcal{N}(\bm{0},\bm{I})$ with a parameterized noise prediction network $\epsilon_{\phi}(\bm{x}_{t},t)$ to predict the noise added to a clean data $\bm{x}_{0}$, which is trained by minimizing
@@
Score Distillation Sampling. Score Distillation Sampling (SDS) is proposed to supervise synthesized novel views by distilling a pretrained text-to-image diffusion model [45] for text-to-3D generation [41]. As the diffusion model is trained on 2D datasets, the SDS loss is referred as $\mathcal{L}_{\textnormal{SDS}}^{2D}$, and its gradient $\nabla_{\theta}\mathcal{L}_{\textnormal{SDS}}^{2D}$ is formulated as:
@@
A pivotal work, Zero-1-to-3 [28] finetunes a variant of Stable Diffusion [44] model on the Objaverse [8] dataset, resulting in a view-conditioned diffusion model $\phi$, which could synthesize novel view images at any arbitrary viewpoint $\bm{c}$ given a reference image $\bm{\hat{I}}$. Hence, $\phi$ offers a strong 3D prior, based on which could derive a 3D-aware SDS loss $\mathcal{L}_{\textnormal{SDS}}^{3D}$. Its gradient $\nabla_{\theta}\mathcal{L}_{\textnormal{SDS}}^{3D}$ is formulated as:
@@ 4 Method
In this section, we propose GeneMAN, a generalizable image-to-3D human pipeline aimed for reconstructing high-quality 3D humans from in-the-wild images. By leveraging GeneMAN prior models, which provide human-specific geometric and texture priors, we achieve high-fidelity 3D human reconstruction with varying body proportions, diverse poses, clothing, and personal belongings. The success of GeneMAN prior models is closely tied to the multi-source human dataset we constructed. We first introduce the construction of a multi-source human dataset and GeneMAN prior models in Sec. 4.1 and Sec. 4.2, respectively. Then, we illustrate the details of our GeneMAN framework, including the Geometry Initialization &\&& Sculpting (Sec. 4.3), as well as the Multi-Space Texture Refinement (Sec. 4.4). The overview of GeneMAN framework is shown in Fig. 2.
Figure 3: Geometry Initialization &\&& Sculpting. During the geometry reconstruction stage, we initialize a template-free geometry using NeRF [36], incorporating GeneMAN 2D and 3D priors with SDS losses. Alongside diffusion-based guidance, a reference loss ensures alignment with the input image. We then convert NeRF into DMTet [52] for high-resolution refinement, guided by pretrained human-specific normal- and depth-adapted diffusion models [17].
@@ 4.1 Multi-Source Human Dataset
The success of LRM [16] has demonstrated that training on multi-source datasets, including the 3D synthetic dataset Objaverse [8] and the video dataset MVImgNet [67], significantly enhances the generalization ability and reconstruction quality of models. However, existing 3D human reconstruction methods typically rely on scarce 3D scanned human data for training, which restricts their ability to generalize to in-the-wild images, particularly human portraits with diverse poses, clothing, and various body proportions. To address this challenge, we construct a large-scale, multi-source human dataset by collecting diverse human data from 3D scans, multi-view videos, single photos, and synthetically generated human data, as illustrated in Fig. 2.
Our 3D scanned human data is aggregated from the commercial dataset RenderPeople [1], alongside several open-source datasets: CustomHumans [12], HuMMan [6], THuman2.0 [66], THuman3.0 [57] and X-Humans [51]. Additionally, we enrich the dataset by integrating human-specific data filtered from Objaverse [8]. For multi-view human videos, we leverage datasets such as DNA-Rendering [7], ZJU-Mocap [40], AIST++ [22], Neural Actor [27] and Actors-HQ [20]. In terms of 2D human imagery, we select data from DeepFashion [29] and LAION-5B [49] to ensure comprehensive coverage of diverse human appearances. Furthermore, we employ data augmentation strategies to synthesize additional data. Specifically, we utilize ControlNet-based [68] image synthesis to generate multi-view human data with diverse clothing options through various prompts, enabling the creation of human images featuring different garments. To account for varying body proportions, we apply image cropping to preprocessed multi-view human renderings, expanding our dataset to include instances with diverse body proportions. In total, our dataset comprises over $50K$ multi-view instances. Further details of the dataset are provided in the supplementary materials.

### v3

@@ Abstract
Given a single in-the-wild human photo, it remains a challenging task to reconstruct a high-fidelity 3D human model. Existing methods face difficulties including a) the varying body proportions captured by in-the-wild human images; b) diverse personal belongings within the shot; and c) ambiguities in human postures and inconsistency in human textures. In addition, the scarcity of high-quality human data intensifies the challenge. To address these problems, we propose a Generalizable image-to-3D huMAN reconstruction framework, dubbed GeneMAN, building upon a comprehensive multi-source collection of high-quality human data, including 3D scans, multi-view videos, single photos, and our generated synthetic human data. GeneMAN encompasses three key modules. 1) Without relying on parametric human models (e.g., SMPL), GeneMAN first trains a human-specific text-to-image diffusion model and a view-conditioned diffusion model, serving as GeneMAN 2D human prior and 3D human prior for reconstruction, respectively. 2) With the help of the pretrained human prior models, the Geometry Initialization-&\&-Sculpting pipeline is leveraged to recover high-quality 3D human geometry given a single image. 3) To achieve high-fidelity 3D human textures, GeneMAN employs the Multi-Space Texture Refinement pipeline, consecutively refining textures in the latent and the pixel spaces. Extensive experimental results demonstrate that GeneMAN could generate high-quality 3D human models from a single image input, outperforming prior state-of-the-art methods. Notably, GeneMAN could reveal much better generalizability in dealing with in-the-wild images, often yielding high-quality 3D human models in natural poses with common items, regardless of the body proportions in the input images.
@@ 1 Introduction
Creating high-quality 3D human models is crucial in various real applications, including VR/AR, telepresence, digital human interfaces, film, and 3D game production. Traditional methods Balan et al. (2007); Liang and Lin (2019); Wu et al. (2020) usually utilize a dense camera array to capture synchronized posed multi-view images for human reconstruction, which typically involves complicated and time-consuming processes. Towards efficient 3D human reconstruction, many approaches Saito et al. (2019); Song et al. (2023); Han et al. (2023); Saito et al. (2020); Huang et al. (2024b); Weng et al. (2024); Zhang et al. (2024c); Ho et al. (2024) delve into the challenge of reconstructing 3D human models from a single image, which, however, remains an ill-posed problem due to the absence of comprehensive 3D human observation data.
@@
To facilitate 3D human reconstruction, previous methods Zheng et al. (2021); Xiu et al. (2022, 2023); Huang et al. (2024b); Zhang et al. (2024c); Ho et al. (2024) often employ parametric human models, e.g., SMPL Loper et al. (2023) or SMPL-X Pavlakos et al. (2019a) as 3D human geometry prior. Nevertheless, these parametric models fail to capture 3D clothing details, which impedes the accurate reconstruction of 3D human figures, particularly when individuals are adorned in loose garments. Recently, image-to-3D human reconstruction methods Ho et al. (2024); Huang et al. (2024b); Zhang et al. (2024c) integrate pretrained text-to-image diffusion models Rombach et al. (2022); Saharia et al. (2022b) to incorporate 2D priors into the reconstruction process. Despite advances in generalizability, single-view 3D human reconstruction has not been fully resolved, especially when dealing with in-the-wild images. The primary challenges are shown in Fig. 1: (a) Varying Body Proportions: many portraits are captured with varying body proportions, such as full-body, half-body, and headshots, while existing methods primarily focus on full-body reconstructions; (b) Human with Objects: in everyday photography, it is common to capture people holding objects, standing on items, or wearing various accessories, which could greatly impact the reconstruction quality. (c) Human Reconstruction with Natural Pose and Textures: due to the absence of a broadly applicable human-specific geometry and texture model, existing methods struggle to reconstruct credible geometry and consistent texture from real-world images. Additionally, the scarcity of high-quality human body data exacerbates the difficulty of tackling the problem.
@@
In this paper, we propose GeneMAN, a Generalizable image-to-3D huMAN reconstruction framework for high-fidelity 3D human reconstruction from a single image. To enhance generalizability, we first collect a comprehensive, multi-source training dataset of high-quality, multi-modal human data, including 3D scans, multi-view videos, single images, and our augmented human data. Based on the multi-source human data collection, human-specific prior models, including a text-to-image diffusion model as the 2D prior and a view-conditioned diffusion model as the 3D prior, have been trained, which could provide more generalizable human priors compared to traditional human parametric models. Leveraging the pretrained human prior models, GeneMAN generates 3D human models through two primary stages. 1) 3D Geometry Reconstruction. To create precise and intricate 3D human geometry, GeneMAN employs an Initialization-&\&-Sculpting strategy, which initially predicts a coarse human geometry using NeRF Mildenhall et al. (2021), followed by a sculpting process for adding geometry details. 2) 3D Texture Generation. Utilizing the refined human geometry, GeneMAN implements a multi-space texture refinement pipeline to produce high-fidelity, consistent 3D textures. Initially, coarse textures are created through multi-view texturing and iteratively refined in the latent space. Subsequently, detailed 3D textures are achieved via pixel space texture refinement, optimizing the UV maps with a 2D prior-based ControlNet Zhang et al. (2023). Extensive experiments demonstrate that GeneMAN surpasses existing state-of-the-art (SoTA) methods, showcasing strong generalization capability and high generation quality. We would like to highlight that GeneMAN could faithfully reconstruct 3D human models with diverse clothing, complex poses, and different personal belongings, given a single in-the-wild image with varying body proportions. Our contributions are summarized as follows:
@@
  • A few effective 3D human reconstruction modules, such as Geometry Initialization &\& Sculpting, and Multi-Space Texture Refinement, have been proposed, facilitating template-free 3D human geometry modeling and view-consistency texturing.
@@ 2 Related Work
Text-to-3D Generation The rapid advancements in 2D image generation have also greatly accelerated progress in text-to-3D generation. DreamFusion Poole et al. (2022) and Magic3D Lin et al. (2023) have demonstrated that utilizing pretrained 2D text-to-image diffusion models Saharia et al. (2022a); Rombach et al. (2022) as guidance can greatly enhance the optimization of 3D representations through Score Distillation Sampling (SDS). Trained on large-scale 2D datasets, these text-to-image models possess excellent capabilities in providing comprehensive 2D knowledge and hallucinating unseen scenes based on specific prompts. Nevertheless, as these models are limited to 2D knowledge, the aforementioned text-to-3D methods are susceptible to multi-view consistency issues. MVDream Shi et al. (2023) improves 3D consistency by training a multi-view diffusion model that offers a robust multi-view prior. However, these methods are not well-suited for direct image-to-3D reconstruction tasks because images cannot be accurately captured through textual descriptions, leading to inconsistencies in color and texture across the generated assets.
Singe Image-to-3D Reconstruction Single image-to-3D reconstruction has greatly benefited from the thriving advancement of diffusion models. Zero-1-to-3 Liu et al. (2023) trains a view-dependent diffusion model on Objaverse Deitke et al. (2023) by explicitly incorporating the camera parameters into the diffusion process. It enables zero-shot image-conditioned novel view synthesis and facilitates general image-to-3D reconstruction with consistent geometric view. Magic123 Qian et al. (2023) and DreamCraft3D Sun et al. (2023) integrate the 3D prior from Zero-1-to-3 Liu et al. (2023) with 2D prior from text-to-image diffusion model, employing a coarse-to-fine two-stage optimization to achieve high-quality 3D reconstruction. LRM Hong et al. (2023) leverages both synthetic data from Objaverse Deitke et al. (2023) and multi-view data from MVImageNet Yu et al. (2023) to train a transformer-based 3D architecture, enhancing its generalizability for 3D reconstruction. However, these general image-to-3D approaches often yield poor results in human reconstruction, resulting in inaccurate geometry of human bodies and the neglect of intricate details, such as facial features and clothing. This limitation arises from the lack of human-specific priors.
@@
Template-based 3D Human Reconstruction. Template-based approaches rely on human parametric models Loper et al. (2023); Pavlakos et al. (2019b); Xu et al. (2020) to guide 3D reconstruction Zheng et al. (2021); Xiu et al. (2022, 2023); Huang et al. (2020); He et al. (2021); Zhang et al. (2024b, c); Ho et al. (2024). PaMIR Zheng et al. (2021) and ICON Xiu et al. (2022) leverage SMPL-derived features to facilitate implicit surface regression, while ECON Xiu et al. (2023) employs explicit representations, predicting normals and depths to generate 2.5D surfaces. GTA Zhang et al. (2024b) utilizes transformers with learnable embeddings to translate image features into 3D triplane features. Several methods Huang et al. (2024b); Ho et al. (2024); Zhang et al. (2024c) exploit the generative capabilities of pretrained diffusion models Rombach et al. (2022) for 3D human reconstruction. SiTH Ho et al. (2024) finetunes an image-conditioned diffusion model to hallucinate back views, while SIFU Zhang et al. (2024c) and TeCH Huang et al. (2024b) utilize diffusion-based geometry and texture refinement to improve reconstruction quality. Recently, IDOL Zhuang et al. (2024) predicts Gaussian-based representation Kerbl et al. (2023) of 3D human to enable fast reconstruction. However, template-based methods are highly reliant on precise human parametric models, which are frequently unattainable for in-the-wild human images.
@@
Template-Free 3D Human Reconstruction. Without parameterized human models, PIFu Saito et al. (2019) introduces a novel approach by extracting pixel-aligned features to build neural fields. PIFuHD Saito et al. (2020) improves this with high-resolution normal guidance, while PHOHRUM Alldieck et al. (2022) jointly predicts 3D geometry, surface albedo, and shading. Recently, HumanSGD AlBahar et al. (2023) leverages pretrained 2D diffusion models Rombach et al. (2022) as a human appearance prior. HumanLRM Weng et al. (2024) trains a diffusion-guided feed-forward model to predict the implicit field of a human without template priors. Nonetheless, they still encounter challenges, such as unrealistic and inconsistent textures, and inferior geometric details, largely due to insufficient priors and overly simplistic model designs.
@@ 3 Preliminaries
Diffusion Model. Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2020) learn to transform samples from a tractable noise distribution towards a data distribution. They comprise a forward process $\{q_{t}\}_{t\in[0,1]}$, which progressively adds random noise to a data point $\bm{x}_{0}\sim q_{0}(\bm{x}_{0})$, and a reverse process $\{p_{t}\}_{t\in[0,1]}$, which gradually recovers clean data from noise. The forward process is defined by a conditional Gaussian distribution: $q_{t}(\bm{x}_{t}|\bm{x}_{0}):=\mathcal{N}(\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$, where $\alpha_{t},\sigma_{t}>0$ are time-dependent coefficients. The reverse process is defined by denoising from $p_{1}(\bm{x}_{1})\coloneqq\mathcal{N}(\bm{0},\bm{I})$ with a parameterized noise prediction network $\epsilon_{\phi}(\bm{x}_{t},t)$ to predict the noise added to a clean data $\bm{x}_{0}$, which is trained by minimizing
@@
Score Distillation Sampling. Score Distillation Sampling (SDS) is proposed to supervise synthesized novel views by distilling a pretrained text-to-image diffusion model Saharia et al. (2022a) for text-to-3D generation Poole et al. (2022). As the diffusion model is trained on 2D datasets, the SDS loss is referred as $\mathcal{L}_{\textnormal{SDS}}^{2D}$, and its gradient $\nabla_{\theta}\mathcal{L}_{\textnormal{SDS}}^{2D}$ is formulated as:
@@
A pivotal work, Zero-1-to-3 Liu et al. (2023) finetunes a variant of Stable Diffusion Rombach et al. (2022) model on the Objaverse Deitke et al. (2023) dataset, resulting in a view-conditioned diffusion model $\phi$, which could synthesize novel view images at any arbitrary viewpoint $\bm{c}$ given a reference image $\bm{\hat{I}}$. Hence, $\phi$ offers a strong 3D prior, based on which could derive a 3D-aware SDS loss $\mathcal{L}_{\textnormal{SDS}}^{3D}$. Its gradient $\nabla_{\theta}\mathcal{L}_{\textnormal{SDS}}^{3D}$ is formulated as:
@@ 4 Method
In this section, we propose GeneMAN, a generalizable image-to-3D human pipeline aimed for reconstructing high-quality 3D humans from in-the-wild images. By leveraging GeneMAN prior models, which provide human-specific geometric and texture priors, we achieve high-fidelity 3D human reconstruction with varying body proportions, diverse poses, clothing, and personal belongings. The success of GeneMAN prior models is closely tied to the multi-source human dataset we constructed. We first introduce the construction of a multi-source human dataset and GeneMAN prior models in Sec. 4.1 and Sec. 4.2, respectively. Then, we illustrate the details of our GeneMAN framework, including the Geometry Initialization &\& Sculpting (Sec. 4.3), as well as the Multi-Space Texture Refinement (Sec. 4.4). The overview of GeneMAN framework is shown in Fig. 2.
@@ 4.1 Multi-Source Human Dataset
The success of LRM Hong et al. (2023) has demonstrated that training on multi-source datasets, including the 3D synthetic dataset Objaverse Deitke et al. (2023) and the video dataset MVImgNet Yu et al. (2023), significantly enhances the generalization ability and reconstruction quality of models. However, existing 3D human reconstruction methods typically rely on scarce 3D scanned human data 70; 45 or only video data Zheng et al. (2024); Xiong et al. (2024); Cheng et al. (2023) for training, which restricts their ability to generalize to in-the-wild images, particularly human portraits with diverse poses, clothing, and various body proportions. To address this challenge, we construct a large-scale, multi-source human dataset by collecting diverse human data from 3D scans, multi-view videos, single photos, and synthetically generated human data, as illustrated in Fig. 2.
Our 3D scanned human data is aggregated from the commercial dataset RenderPeople 45, alongside several open-source datasets: CustomHumans Ho et al. (2023), HuMMan Cai et al. (2022), THuman2.0 Yu et al. (2021), THuman3.0 Su et al. (2022) and X-Humans Shen et al. (2023). Additionally, we enrich the dataset by integrating human-specific data filtered from Objaverse Deitke et al. (2023). For multi-view human videos, we leverage datasets such as DNA-Rendering Cheng et al. (2023), ZJU-Mocap Peng et al. (2021), AIST++ Li et al. (2021), Neural Actor Liu et al. (2021) and Actors-HQ Işık et al. (2023). In terms of 2D human imagery, we select data from DeepFashion Liu et al. (2016) and LAION-5B Schuhmann et al. (2022) to ensure comprehensive coverage of diverse human appearances. Furthermore, we employ data augmentation strategies to synthesize additional data. Specifically, we utilize ControlNet-based Zhang et al. (2023) image synthesis to generate multi-view human data with diverse clothing options through various prompts, enabling the creation of human images featuring different garments. To account for varying body proportions, we apply image cropping to preprocessed multi-view human renderings, expanding our dataset to include instances with diverse body proportions. In total, our dataset comprises over $50K$ multi-view instances. Further details of the dataset are provided in the supplementary materials.